I am just drawing a blank on what the 2 equations would be for this problem. Once I get them. I can solve the problem.
Simone can build a treehouse in 15 hours. Working together Simone and Tyrese can build a treehouse in 7 hours. How many hours would Tyrese take to build the treehouse working alone?
I have S=15
S+T=7 but I know this isn't correct.
Please help

Found 2 solutions by stanbon, scott8148:
Answer by stanbon(48568)

Simone can build a treehouse in 15 hours. Working together Simone and Tyrese can build a treehouse in 7 hours. How many hours would Tyrese take to build the treehouse working alone?
--------------------
Simone DATA:
Time = 15 hrs/job ; Rate = 1/15 job/hr
------------------------
Together DATA:
Time = 7 hrs/job ; Rate = 1/7 job/hr
------------------------
Tyrese DATA:
Time = x hrs/job ; Rate = 1/x job/hr
-------------------------
EQUATION:
rate + rate = rate together
1/15 + 1/x = 1/7
Multiply thru by 105x to get:
7x + 105 = 15x
8x = 105
x = 13.125 hrs (time it takes Tyrese to build the tree house alone)
=================
Cheers,
Stan H.

Answer by scott8148(5894)

with work problems, the trick is to see what fraction of the job is done by each person

if Simone can build a treehouse in 15 hours, then in 7 hours she builds 7/15 of the treehouse

this means Tyrese builds 8/15 of the tree house in 7 hours ... so 7 is to 8/15 as x is to 15/15 (or 1 whole treehouse)